OpenStudy (anonymous):

HELP PLEASE! y = -3/2x + 4 Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. (the / in -3/2 is a fraction bar)

OpenStudy (austinl):

Ok.... my webpage crashed. Let me type that ALL back out again.... :/

OpenStudy (anonymous):

ok thanks

OpenStudy (austinl):

\(\large{y=-\frac{3}{2}x+4}\) 4 is your y intercept, so you would plot a point at (0,4) Your slope is \(-\dfrac{3}{2}\rightarrow \dfrac{rise}{run}\) That means that we would go down three from our original point, and right two. (2,1) Then you draw a line through the two plotted points.

OpenStudy (anonymous):

you are amazing!!

OpenStudy (austinl):

No problemo :)

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

one more question please?

OpenStudy (anonymous):

–3x + 2y = 8 Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.

OpenStudy (austinl):

This one you need to solve for y.

OpenStudy (austinl):

\(\large{-3x+2y=8}\) \(\large{2y=3x+8}\) \(\large{y=\frac{3}{2}x+8}\) Can you do it from there given what I just told you?

OpenStudy (anonymous):

well the y is already isloated

OpenStudy (austinl):

Correct, I was referring to this in comparison to the previous problem. The only difference is the slope. Instead of going down three, we go up three.

OpenStudy (anonymous):

how do i write this in a complete sentence?

OpenStudy (anonymous):

Plot[4 - (3 x)/2, {x, -5.3, 5.3}]

OpenStudy (austinl):

It is literally exactly the same as the first problem..... just instead of going down three, we go up three.

OpenStudy (anonymous):

so that means that we would go up three from our original point and right two.

OpenStudy (anonymous):

but i still dont understand how i would write this in a complete sentence with the plots. but what are the plots?